OpenStudy (anonymous):

hey guys I need help in this question Determine the 8th term of the following arithmetic sequence. -7, 2, 11, 20

OpenStudy (nirmalnema):

nth term=a+(n-1)d a=-7 d=9 n=8 -7+(8-1)9 -7+63 =56